OpenStudy (anonymous):

Solve the following formula for W: x=2W^2/Z

OpenStudy (zehanz):

So you have:\[x=\frac{ 2W^2 }{ Z }\]It already help if you had:\[2W^2=...\] If you have trouble doing that, imagine you had\[6=\frac{ 18 }{ 3 }\]The you would know in an instant what to do to write it as:\[18=...\]

OpenStudy (anonymous):

 I have this so far though?

OpenStudy (zehanz):

Looks good! So now you have to get from w² to w...

OpenStudy (anonymous):

Thats the part i dont know how to do

OpenStudy (anonymous):

 IS IT RIGHT?

OpenStudy (zehanz):

You mean take the root of xz/2? Then you are right!

OpenStudy (anonymous):

Or do i keep the square\[\sqrt{XZ/2}^2\]d? \[\sqrt{XZ/2} =W\]

OpenStudy (anonymous):

WAIT SO WHATS IS IT? D: i HAVE \[\sqrt{xz/2}\] I have this as an answer but i dont know if its right

OpenStudy (zehanz):

Strictly spoken, there are two possibilities: \[W=\pm \sqrt{\frac{ xz }{ 2 }}\]To understand this, take a look at this: w²=16. What is w? Well, because 4²=16, w=4. But: (-4)²=16 as well, so w=-4 is also correct. There are two possibilities!

OpenStudy (anonymous):

Thank you so much!!!!!!!!!